OpenStudy (anonymous):

Solve the equation then check your solution. 3.7(-83x-3) = 3.7(7x-30) I keep getting x=1/3 but when I go to check it, it's wrong.

OpenStudy (anonymous):

3.7(-83x-3)=3.7(7x-30) do each side separately

OpenStudy (anonymous):

3.7(-83x-3) = -307.1x-11.1 3.7(7x-30) = 25.9x-111

OpenStudy (anonymous):

-307.1x-11.1 = 25.9x - 111

OpenStudy (anonymous):

-307.1x-11.1 = 25.9x - 111 +11.1 +11.1 -307.1x = 25.9x - 99.9 -25.9x -25.9x -333x = - 99.9 divide both sides by -333 x = (3/10)

OpenStudy (anonymous):

do you understand?